OpenStudy (anonymous):

Solve the equation. |10x – 14| - 5 = 156

OpenStudy (crashonce):

move -5 and -14 to RHS

OpenStudy (crashonce):

10x = 137, x = 13.7

OpenStudy (shadowlegendx):

@CrashOnce could you inform the individual and I what RHS stands for?

OpenStudy (crashonce):

RHS stands for Right Hand Side LHS stands for Left Hand Side Australian style c:

OpenStudy (anonymous):

well you were wrong

OpenStudy (aum):

|10x – 14| - 5 = 156 add 5 to both sides: |10x – 14| = 161 If |a| = b then it implies: a = b or a = -b Therefore, |10x – 14| = 161 implies: 10x - 14 = 161 or 10x - 14 = -161 10x - 14 = 161 10x = 161+14 = 175 x = 17.5 or 10x - 14 = -161 10x = -161 + 14 = -147 x = -14.7 x = -14.7 or 17.5